# Match, Compare, or Select? An Investigation of Large Language Models for Entity Matching

### v1

@@ 4.3 LLM as a Selector
In this section, we propose a selecting strategy that uses an LLM to select the matching record of a given record from a list of candidate matches. Specifically, given an entity record $r$ and its candidate matches $R_{c}=\{r_{1},r_{2},\ldots,r_{n}\}$, this strategy directly selects the match of record $r$ from $R_{c}$ by interfacing LLMs with an appropriate selection prompt $\operatorname{P}_{s}$, as shown in Figure 1(c):
$\operatorname{LLM}_{s}\colon\left\{(r,R_{c})\right\}\xrightarrow{\operatorname{P}_{s}}\left\{1,2,\ldots,n\right\}$
@@
In this way, LLMs can be explicitly required to identify only one match per record $r$ in the candidate matches $R_{c}$. Furthermore, feeding LLMs all candidate matches in the same context at a time allows LLMs to make better decisions by considering interactions between candidate records. For example, if “SanDisk Cruzer Glide”, “SanDisk Cruzer Force”, and “SanDisk Extreme Pro” appear in different records of the same context, it is easier for LLMs to recognize that they are different model names of SanDisk flash drives and make the correct choice for the actual match.
One challenge in applying the selecting strategy to LLM-based entity matching is that there is often no actual match of record $r$ in candidate matches $R_{c}$, because entity matching is an extremely imbalanced task. A trivial solution to this challenge could be to perform a pairwise matching after the selection, which would undermine the advantages of the selecting strategy. Another method could be to add “none of the above” as an additional option to allow LLMs to refuse to select any record from the candidate matches. Suppose the selection prompt with the “none of the above” option is $\operatorname{P}_{s_{N}}$, this method can be formulated as:
$\operatorname{LLM}_{s_{N}}\colon\left\{(r,R_{c})\right\}\xrightarrow{\operatorname{P}_{s_{N}}}\left\{0,1,2,\ldots,n\right\}$
@@
However, the selecting strategy relies heavily on the capabilities of LLMs for fine-grained understanding and implicit ranking in long contexts. Our experimental results (§ 5.2) show that the current LLMs suffer from position bias, with the selection accuracy decreasing significantly as the position of the matching record in the candidate list increases. In practice, the recall-oriented blocking step often generates a considerable number of candidate matches for each record, exceeding the context length that LLMs can effectively handle. It is a challenge to mitigate the position bias and the long context requirement for the selecting strategy.
4.4 ComEM Framework
@@
# Prompt
Tokens
Matching $\mathcal{O}(n)$ $\mathcal{O}(1)$
Comparing $\mathcal{O}(n^{2})$ or $\mathcal{O}(n)$ $\mathcal{O}(1)$
Selecting $\mathcal{O}(1)$ $\mathcal{O}(n)$
@@
Table 1: Comparison of different strategies. “” means that the matching strategy can only calibrate similarity scores if the generation probability is available.
@@
Given the advantages and shortcomings of different strategies, we design a compositional entity matching (ComEM) framework. This ComEM framework addresses various challenges of entity matching by integrating the advantages of different strategies and LLMs. Table 1 shows a comparison of these strategies. The matching and comparing strategies are applicable for local ranking, while the selecting strategy is suitable for fine-grained identification. Therefore, as shown in Figure 2, we first utilize a medium-sized LLM to rank and filter the candidate matches $R_{c}$ of record $r$ with the matching or comparing strategy. We then utilize an LLM to identify the match of record $r$ from only top $k$ candidates with the selecting strategy.
@@
Our ComEM framework integrates the advantages of different strategies through a filtering then identifying pipeline. It first utilizes the local matching or comparing strategy to rank candidate matches for preliminary screening, which can effectively mitigate the position bias and the long context requirement of the selecting strategy. It then utilizes the global selecting strategy to incorporate record interactions for fine-grained optimization, which can effectively mitigate the consistency ignorance of the matching strategy. Thus, ComEM is able to strike a balance between entity matching requirements and current LLM capabilities, achieving significant performance improvements.
@@
By integrating LLMs of different sizes, our ComEM framework can also effectively reduce the cost of LLM invocations for entity matching. It is expensive to use commercial LLMs directly in practice because entity matching is a computationally intensive task. Our ComEM framework delegates a significant part of the computation to medium-sized LLMs. Experimental results reveal that the ranking process can be performed well by using open-source medium-sized (3B~11B) LLMs. Therefore, the proper integration of LLMs not only improves the performance of entity matching but also reduces the cost in practice.
Figure 2: Illustration of our ComEM framework.
@@
Abt-Buy 44.23 29.32 90.00 0.25 64.19 60.00 69.00 0.64 67.94 54.94 89.00 0.09
Amazon-Google 35.47 25.00 61.00 0.24 45.88 55.71 39.00 0.62 53.79 43.29 71.00 0.09
DBLP-ACM 67.35 51.03 99.00 0.35 77.36 73.21 82.00 0.94 70.04 54.80 97.00 0.15
DBLP-Scholar 66.13 55.41 82.00 0.33 62.65 78.79 52.00 0.84 67.40 53.18 92.00 0.12
IMDB-TMDB 65.82 89.66 52.00 0.27 44.12 83.33 30.00 0.70 85.84 78.99 94.00 0.11
IMDB-TVDB 57.59 60.44 55.00 0.24 52.78 86.36 38.00 0.60 72.73 69.72 76.00 0.08
TMDB-TVDB 69.43 55.76 92.00 0.24 80.19 75.89 85.00 0.61 71.60 58.60 92.00 0.08
Walmart-Amazon 41.84 27.16 91.00 0.39 62.56 64.21 61.00 1.03 65.64 53.46 85.00 0.17
Mean 55.98 49.22 77.75 0.29 61.22 72.19 57.00 0.75 69.37 58.37 87.00 0.11
Table 2: Overall performance and cost of different strategies. We bold the best F1 score.
5 Experiments
In this section, we conduct thorough experiments to investigate three strategies for entity matching with LLMs. First, we perform the analysis of different strategies (§ 5.2). Next, we examine the effect of different LLMs on these strategies (§ 5.3). Finally, we verify our ComEM framework that is proposed based on investigation findings (§ 5.4).
@@ 5.1 Experimental Setup
Dataset Domain # D1 # D2 # Attr # Pos
Abt-Buy Product 1076 1076 3 1076
Amazon-Google Software 1354 3039 4 1103
DBLP-ACM Citation 2616 2294 4 2224
DBLP-Scholar Citation 2516 61353 4 2308
IMDB-TMDB Movie 5118 6056 5 1968
IMDB-TVDB Movie 5118 7810 4 1072
TMDB-TVDB Movie 6056 7810 6 1095
Walmart-Amazon Electronics 2554 22074 6 853
Table 3: Statistics of experiment datasets.
Datasets.
@@
We focused on the common record linkage that has many open-access datasets. We used 8 ER datasets collected by pyJedAI Nikoletos et al. (2022) because pairwise EM benchmarks Mudgal et al. (2018) for deep learning-based methods involve extensive negative pair downsampling, which conceals real challenges of EM Papadakis et al. (2023). For each dataset with two record collections D1 and D2, we retrieved 10 candidate matches from D2 for each record in D1 using the SOTA blocking method Sparkly Paulsen et al. (2023). The recall of Sparkly on all datasets ranges from 86.57% to 99.96%, demonstrating its effectiveness in retrieving candidate matches. This allows us to investigate in real-world scenarios. Table 3 summarizes the statistics of these datasets.
Evaluation Metrics.
@@
Consistent with prior studies, we report F1, Precision, and Recall on record pairs as evaluation metrics for entity matching. We also report the cost (\$) of LLM invocations [^1] Inference cost is estimated based on the hourly price of the cloud NVIDIA A40..
Implementation Details.
We used ChatGPT (gpt-3.5-turbo-0613) as the main LLM for strategy analysis. For the comparing strategy, we match the top 1 candidate after ranking. We also examined the effect of 9 open-source LLMs on these strategies, including Llama2-7/13/70B-Chat Touvron et al. (2023), Mistral-7B Jiang et al. (2023), Mixtral-8x7B Jiang et al. (2024), Baichuan2-13B-Chat Yang et al. (2023), Vicuna1.5-13B Zheng et al. (2023), SOLAR-10.7B Kim et al. (2023), and Flan-T5-11B Chung et al. (2022). The generation temperature was set to 0 for reproducibility. We sampled 200 records for each dataset, half of which had matches, and formed 2000 pairs of records with their candidate matches. In ComEM, we used Flan-T5-11B to rank candidates with the matching strategy and kept the top 4 candidates for selection.
@@ 5.2 Analysis of Strategies
We first compare the performance, cost, and bias of the strategies, with the following findings.
Figure 3: Comparison of different strategies.
Figure 4: Effect of open-source LLMs on different strategies. Star pentagons represent the results of ChatGPT.
Figure 5: F1 score w.r.t. matched record positions.
Finding 1. Incorporating record interactions is essential for entity matching with LLMs. As shown in Table 2, entity matching performance increases as the level of record interaction increases. The comparing strategy outperforms the independent matching strategy by an average of 5.24% F1 score, and the selecting strategy further improves the F1 score by up to 8.15% over the comparing strategy. The selection strategy is optimal in 6 out of 8 datasets, and suboptimal in the remaining two, fully demonstrating the importance of incorporating record interactions. To further verify that these improvements are brought by the strategy, we also perform 6-shot matching, ensuring that the number of records is consistent with the selecting strategy. As shown in Figure 3, the selecting strategy still has a significant F1 advantage over 6-shot matching. These results highlight the importance of incorporating record interactions and open new avenues for entity matching with LLMs.
Finding 2. The selecting strategy is the most cost-effective for entity matching with LLMs. Monetary cost is also an important factor when interfacing LLMs for EM in practice, as it is computationally intensive. As shown in Table 2, the selecting strategy costs less than half of the matching strategy. Although the complexity of the number of prompt tokens for these two strategies is $\mathcal{O}(n)$, the number of times a record is repeatedly prompted to the LLM, as well as the repetition of task instructions, can significantly affect the actual cost. The comparing strategy, however, considers two candidate matches at a time and interfaces the LLM twice, making its cost more than twice that of the matching strategy. Therefore, the selecting strategy stands out for its effectiveness and efficiency.
Finding 3. Strategies that incorporate multiple records suffer from the position bias of LLMs. As shown in Figure 5, the performance of the comparing and selecting strategies decreases significantly as the position of the matching records moves down in the candidate list. For the comparing strategy optimized with bubbling sort, the matching records cannot be ranked at the top if there is any incorrect comparison. LLMs struggle with comparing records, as we will further see in the next section. The selecting strategy also drops nearly 10% in F1, probably due to the limited long context understanding of the LLM. As a result, the position bias of LLMs restricts the performance and generality of the comparing and selecting strategies.
@@ 5.3 Effect of LLMs
Finding 4. There is no single LLM that is uniformly dominant across all strategies. Figure 4 shows the efficacy of open-source LLMs on these strategies, with detailed results in Appendix A. We can see that most LLMs perform well in the selecting strategy, but may perform quite poorly in the matching and comparing strategies. This consistent performance of LLMs in the selecting strategy once again confirms the effectiveness of incorporating global information in this way. For the current matching strategy, however, common chat LLMs, especially the Llama2 series and Vicuna1.5, lack the capability to identify matches and produce a large number of false positives. This suggests that LLMs that supervised fine-tuning only on conversation data may not be directly applicable to specific tasks. The situation is even worse for the comparing strategy. Almost all LLMs, even the powerful Mixtral-8x7B, are unable to find more similar records from two candidate matches. These LLMs either make inconsistent choices between two comparisons that exchange the candidate order, or they always output a specific identifier regardless of the records. Although Flan-T5-11B lags behind ChatGPT by about 10% F1 in the selecting strategy, we find that it performs quite well in the matching and comparing strategies. This instruction-tuned LLM follows instructions better than chat-tuned LLMs and can output only the requested labels instead of long-form responses, making it convenient to utilize label generation probabilities. In conclusion, there is a noticeable variance in the capabilities of different LLMs for a single strategy, and the efficacy of different strategies for a single LLM can also be significantly distinct.
@@
Finding 5. Matching strategy is better for ranking and filtering than comparing strategy. The superiority of Flan-T5 in the matching and comparing strategies leads us to explore the possibility of using it to rank and filter candidate matches for the selecting strategy. As shown in Figure 6, the matching strategy outperforms the comparing strategy under different model parameter sizes, even though the latter performs $\mathcal{O}(n^{2})$ comparisons. Their difference is small on the 3B flan-t5-xl and 11B flan-t5-xxl models, but significant on smaller models. This may be due to the fact that these models are trained on many pairwise tasks, such as natural language inference and question answering, but few triple-wise tasks. Therefore, in terms of effectiveness and efficiency, the matching strategy is more suitable for ranking and filtering candidate matches.
5.4 Compositional Entity Matching
We finally verify the feasibility of our ComEM that integrates multiple strategies and LLMs.
Finding 6. By integrating the advantages of different strategies and LLMs, ComEM can accomplish EM more effectively and cost-efficiently. As shown in Figure 7, compared to the optimal selecting strategy using ChatGPT, ComEM achieves a significant F1 improvement (~8%) while spending less. The filtering and identifying pipeline improves precision considerably without sacrificing high recall of the selecting strategy. These results reveal that integrating multiple strategies can complement single strategies and effectively mitigate the position bias of the selecting strategy in long contexts. However, using a single powerful but costly commercial LLM to complete the entire pipeline obscures the cost efficiency of the selecting strategy. By introducing a medium-sized but good at matching LLM for preliminary filtering, ComEM maintains the performance improvement and even spends less than direct selection. As a result, our ComEM framework underscores the importance of task decomposition and LLM composition, illuminating an effective route for compositional entity matching with LLMs.
Figure 7: Performance and cost of compositional entity matching. “M.” and “S.” denote matching and selecting.
Ablation Study.
We perform an ablation study on the number of candidate records for further identification. As shown in Figure 8, recall increases and precision decreases as the number of retained candidate matches increases. Consistent with Figure 5, four is the sweet spot for the selecting strategy with current ChatGPT, which balances precision and recall to achieve the best F1 score.
Figure 8: Average F1, precision, and recall w.r.t. number of candidate matches ($k$) for further identification. See Appendix B for detailed results on each dataset.

### v2

@@ 3.4 LLM as a Selector
In this section, we introduce a selecting strategy that uses an LLM to select the matching record of a given record from a list of potential matches. Specifically, given an entity record $r$ and its potential matches $R=\{r_{1},r_{2},\ldots,r_{n}\}$, this strategy directly selects the match of record $r$ from $R$ by interfacing LLMs with an appropriate selection prompt, as shown in Figure 1(c):
$\operatorname{LLM}_{s}\colon\left\{(r,R)\right\}\rightarrow\left\{1,2,\ldots,n\right\}$
@@
In this way, LLMs can be explicitly required to identify only one match per record $r$ from the potential matches $R$. Furthermore, feeding LLMs all potential matches in the same context at a time allows LLMs to make better decisions by considering interactions between records. For example, if “SanDisk Cruzer Glide”, “SanDisk Cruzer Force”, and “SanDisk Extreme Pro” appear in different records of the same context, it is easier for LLMs to recognize that these are different model names of SanDisk flash drives and select the actual match.
One challenge in applying the selecting strategy to LLM-based entity matching is that there is often no actual match of record $r$ in potential matches $R$, because entity matching is an imbalanced task. A trivial solution to this challenge could be to perform a pairwise matching after the selection, which would undermine the advantages of the selecting strategy. Another method could be to add “none of the above” as an additional option to allow LLMs to refuse to select any record from the potential matches, which can be formulated as:
$\operatorname{LLM}_{s_{N}}\colon\left\{(r,R)\right\}\rightarrow\left\{0,1,2,\ldots,n\right\}$
@@
However, the selecting strategy relies heavily on the capabilities of LLMs for fine-grained understanding and implicit ranking in long contexts. Our experimental results show that the current LLMs suffer from position bias, with the selection accuracy decreasing significantly as the position of the matching record increases in the candidate list (§ 4.3). In practice, the recall-oriented blocking step often generates a considerable number of potential matches for each record, exceeding the context length that LLMs can effectively reasoning (Levy et al., 2024). Therefore, it is a challenge to mitigate the position bias and the long context requirement for the selecting strategy.
3.5 Compound Entity Matching Framework
@@
# Input
Records
Matching $\mathcal{O}(n)$ $2n$
Comparing $\mathcal{O}(kn)$ $3k(2n-k-1)$
Selecting $\mathcal{O}(1)$ $n+1$
@@
Table 1: Comparison of different strategies. “” means that the matching strategy can only calibrate similarity scores if the generation probability is available. “# Input Records” represents the number of (#) records input to LLMs using different strategies for record $r$ and its $n$ potential matches $R$. $k$ denotes the number of top candidates considered by the comparing strategy.
@@
Based on the advantages and shortcomings of different strategies, we further propose a compound entity matching framework (ComEM). ComEM addresses various challenges in LLM-based entity matching by integrating the advantages of different strategies and LLMs. Table 1 shows a comparison of these strategies. The matching and comparing strategies are applicable for local ranking, while the selecting strategy is suitable for fine-grained identification. Therefore, as shown in Figure 2, we first utilize a medium-sized LLM to rank and filter potential matches $R$ of record $r$ with the matching or comparing strategy. We then utilize an LLM to identify the match of record $r$ from only the top $k$ candidates with the selecting strategy.
@@
Our ComEM framework integrates the advantages of different strategies through a filtering then identifying pipeline. It first utilizes the local matching or comparing strategy to rank potential matches for preliminary screening, which can effectively mitigate the position bias and the long context requirement of the selecting strategy. It then utilizes the global selecting strategy to incorporate record interactions for fine-grained optimization, which can effectively mitigate the consistency ignorance of the matching strategy. Therefore, ComEM is able to strike a balance between entity matching requirements and current LLM capabilities, achieving significant performance improvements.
@@
By integrating LLMs of different sizes, our ComEM framework can also effectively reduce the cost of LLM invocations for entity matching. In practice, direct use of commercial LLMs is expensive because entity matching is a computationally intensive task. ComEM delegates a significant part of the computation to medium-sized LLMs. Experimental results show that the ranking process can be performed well by using open-source medium-sized (3B~11B) LLMs (§ 4.4). As a result, the proper integration of LLMs not only improves the performance of entity matching but also reduces the cost for practical application.
Figure 2: Illustration of ComEM framework. It first filters candidate records by matching or comparing strategies and then identifies the match via selecting strategy.
4 Experiments
In this section, we conduct thorough experiments to investigate three strategies for LLM-based entity matching. First, we present the main experimental results (§ 4.2). Next, we perform the analysis of different strategies (§ 4.3). Finally, we examine the effect of different LLMs on these strategies (§ 4.4).
Dataset Domain # D1 # D2 # Attr # Pos
Abt-Buy (AB) Product 1076 1076 3 1076
Amazon-Google (AG) Software 1354 3039 4 1103
DBLP-ACM (DA) Citation 2616 2294 4 2224
DBLP-Scholar (DS) Citation 2516 61353 4 2308
IMDB-TMDB (IM) Movie 5118 6056 5 1968
IMDB-TVDB (IV) Movie 5118 7810 4 1072
TMDB-TVDB (TT) Movie 6056 7810 6 1095
Walmart-Amazon (WA) Electronics 2554 22074 6 853
Table 2: Statistics of experimental datasets.
Method Metric AB AG DA DS IM IV TT WA Mean Cost
P 71.43 38.00 86.71 71.19 84.44 57.81 67.97 73.49 68.88
R 50.00 75.00 95.67 84.00 85.00 95.00 87.00 40.67 76.54
Sudowoodo F1 58.82 50.45 90.97 77.06 84.72 71.88 76.32 52.36 70.32 1.11
P 40.41 35.54 65.78 64.63 95.08 68.75 65.28 35.62 58.89
R 91.33 59.00 98.67 81.00 58.00 55.00 94.00 88.33 78.17
Matching F1 56.03 44.36 78.93 71.89 72.05 61.11 77.05 50.77 64.02 4.52
P 81.69 65.31 85.60 82.74 96.55 84.82 88.93 71.26 82.11
R 77.33 42.67 69.33 54.33 46.67 31.67 85.67 60.33 58.50
Comparing F1 79.45 51.61 76.61 65.59 62.92 46.12 87.27 65.34 66.86 11.75
P 74.08 58.13 81.34 73.89 89.41 84.07 77.18 72.95 76.38
R 87.67 70.33 97.33 88.67 95.67 82.67 91.33 89.00 87.83
[…]
P 85.67 66.57 86.23 80.48 94.59 86.06 79.94 85.11 83.08
R 89.67 73.00 96.00 89.33 99.00 82.33 90.33 87.67 88.42
@@ 4.1 Experimental Setup
Datasets.  We focused on the common record linkage that has many open-access datasets. Specifically, we used 8 clean-clean ER datasets collected by pyJedAI Nikoletos et al. (2022). Table 2 shows the statistics of these datasets. We adapted them to fit the problem formulation and to support our experiments. For each dataset with two record collections D1 and D2, we applied the SOTA blocking method Sparkly Paulsen et al. (2023) to retrieve 10 potential matches from D2 for each record in D1. The recall@10 of Sparkly on all datasets ranges from 86.57% to 99.96%, demonstrating its effectiveness in retrieving potential matches. In this way, we are able to investigate and evaluate different strategies under the real ER pipeline.
Baseline.  Except for the pairwise matching strategy, we also compare the STOA self-supervised learning method, Sudowoodo Wang et al. (2023), which reduces the need for supervision through contrastive learning and pseudo-labeling.
@@
Evaluation Metrics.  Consistent with prior studies, we report F1, Precision, and Recall on record pairs. We also report the cost (\$) of LLM invocations [^2] The inference or training cost is estimated based on the hourly price of the cloud NVIDIA A40..
Implementation Details. [^3] We have provided the full code, including blocking and sampling in the Supplementary Material for reproducibility. We used ChatGPT (gpt-3.5-turbo-0613) as the main LLM for strategy analysis. We also examined the effect of 8 open-source instruction-tuned LLMs, including Llama-3-8B (AI@Meta, 2024), Qwen2-7B (Bai et al., 2023), Mistral-7B (Jiang et al., 2023), Mixtral-8x7B (Jiang et al., 2024), Flan-T5-XXL (Chung et al., 2022), Flan-UL2 (Tay et al., 2023) and Solar-10.7B (Kim et al., 2023). The specific prompts can be found in Appendix A, with the generation temperature of all LLMs set to 0 for reproducibility. For each dataset, we sampled 400 records from record collection D1 for evaluation, 300 of which had matches, and formed 4000 pairs of records with their potential matches obtained by Sparkly from record collection D2. The remaining record pairs (unsampled records and their potential matches) are used for model training or in-context learning. For Sudowoodo, we used its official implementation [^4] https://github.com/megagonlabs/sudowoodo to train models on 500 pairs. For in-context learning, we select 100 record pairs and follow Peeters and Bizer (2023a) to retrieve 3 positives and 3 negatives as few-shot examples. Since the comparing strategy produces only relative orders, we applied the matching strategy to the top 1 candidate after bubble sort ranking. In ComEM, we used Flan-T5-XL to rank candidates with the matching strategy and kept the top 4 candidates for selection.
4.2 Main Results
Figure 3: Comparison of different strategies.
We first compare the performance and cost of different methods, with the following findings.
Finding 1. Incorporating record interactions is essential for LLM-based entity matching. As shown in Table 3, entity matching performance increases as the level of record interaction increases. The comparing strategy outperforms the independent matching strategy by an average of 2.84% F1 score, and the selecting strategy further improves the F1 score by up to 14.74% over the comparing strategy. The optimal selecting strategy is 11.28% higher in F1 than the SOTA self-supervised learning method. The advantages of the comparing and selecting strategies over the matching strategy are also evident across different LLMs in Figure 5. To further verify that these improvements are due to the strategy, we perform 6-shot matching, ensuring that the number of records is consistent with the selecting strategy. As shown in Figure 3, the selecting strategy still has a significant F1 advantage over 6-shot matching. These results highlight the effectiveness of our proposed strategies and open new avenues for LLM-based entity matching.
Finding 2. By integrating the advantages of different strategies and LLMs, ComEM can accomplish EM more effectively and cost-efficiently. As shown in Table 3 and Figure 3, compared to the optimal selecting strategy using ChatGPT, ComEM achieves 4% F1 improvement while spending less. The filtering and identifying pipeline improves precision considerably (6.7%) without sacrificing high recall of the selecting strategy. These results reveal that integrating multiple strategies can complement single strategies and mitigate the position bias of the selecting strategy in long contexts. However, using a single powerful but costly commercial LLM to complete the entire pipeline obscures the cost efficiency of the selecting strategy. By introducing a medium-sized LLM for preliminary filtering, ComEM improves performance while spending less than direct selection. As a result, ComEM underscores the importance of task decomposition and LLM composition, illuminating an effective route for compound entity matching using LLMs.
Figure 4: F1 score w.r.t. matched record positions.
@@ 4.3 Analysis of Strategies
We then analyze the advantages and shortcomings of different strategies from different perspectives.
Finding 3. The selecting is the most cost-effective strategy for LLM-based entity matching. Monetary cost is also an important factor when interfacing LLMs for EM in practice, as it is computationally intensive. As shown in Table 3, the selecting strategy costs less than half of the matching strategy. This is because the selecting strategy saves $n-1$ times of repeatedly inputting anchor records and task instructions into LLMs. The comparing strategy, however, considers two potential matches at a time and interfaces the LLM twice, making its cost more than twice that of the matching strategy. Therefore, the selecting strategy stands out for its effectiveness and efficiency.
Finding 4. Strategies that incorporate multiple records suffer from the position bias of LLMs. As shown in Figure 4, the performance of the comparing and selecting strategies decreases significantly as the position of the matching records moves down in the candidate list. For the comparing strategy optimized with bubble sort, matching records cannot be ranked at the top if there is any incorrect comparison. The selecting strategy also drops about 10% in F1, probably due to the limited long context understanding of the LLM. Therefore, the position bias of LLMs restricts the performance and generality of the comparing and selecting strategies.
Figure 5: Effect of open-source LLMs on different strategies.
@@ 4.4 Effect of LLMs
Finding 5. There is no single LLM that is uniformly dominant across all strategies. Figure 5 shows the efficacy of proposed strategies for open-source LLMs, with detailed results in Appendix B. We can see that the F1 scores of the matching, comparing, and selecting strategies for different LLMs mostly fall between 50%~70%, 60%~80%, and 70%~80%, respectively. In general, similar to ChatGPT, the comparing strategy is better than the matching strategy, while the selecting strategy is further better than the comparing strategy. The consistent performance between strategies confirms the effectiveness of incorporating record interactions in these ways. Concretely, some chat LLMs such as Llama3-8B and Mistral-7B produce numerous false positives and thus perform poorly with the matching strategy. Nevertheless, they achieve significant improvement and satisfactory performance by comparing and selecting strategies, respectively. Moreover, although Flan-T5-XXL and Flan-UL2 lag behind ChatGPT by about 4% F1 in the selecting strategy, we find that they perform quite well in the matching and comparing strategies. These task-tuned LLMs follow instructions better than chat-tuned LLMs, and can output only the requested labels instead of long-form responses, making it convenient to utilize label generation probabilities. In conclusion, there is a noticeable variance in the capabilities of different LLMs for a single strategy, and the efficacy of different strategies for a single LLM can also be significantly distinct.
@@
Finding 6. Matching strategy is better for ranking and filtering than comparing strategy. The superiority of Flan-T5 in the matching and comparing strategies leads us to explore the possibility of using it to rank and filter potential matches for the selecting strategy. As shown in Figure 6, the matching strategy outperforms the comparing strategy under different model parameter sizes, even though the latter performs $\mathcal{O}(n^{2})$ comparisons. The difference is small on Flan-T5-XL (3B) and Flan-T5-XXL (11B), but significant on smaller models. This may be due to the fact that these models are trained on many pairwise tasks, such as natural language inference and question answering, but few triple-wise tasks. Therefore, in terms of effectiveness and efficiency, the matching strategy is more suitable for ranking and filtering potential matches.
4.5 Ablation Study
We perform an ablation study on the number of candidate records for further identification. As shown in Figure 7, recall increases and precision decreases as the number of retained potential matches increases. Consistent with Figure 4, four is the sweet spot for the selecting strategy with current ChatGPT, which balances precision and recall well.
Figure 7: Average F1, precision, and recall w.r.t. number of candidate retained ($k$) for further selection. See Appendix C for detailed results on each dataset.